OpenStudy (anonymous):

Over which interval does the function decrease the fastest? f(x) = 3√ (-x + 3) A.4 < x < 6 B.0 < x < 2 C.6 < x < 8 D.2 < x < 4

OpenStudy (anonymous):

what level math is this ?

OpenStudy (anonymous):

Nonlinear algebra

OpenStudy (anonymous):

id say its C

OpenStudy (cwrw238):

no - x =6 to 8 would give negative square root

OpenStudy (cwrw238):

the graph would finish at the point (0,3)

OpenStudy (anonymous):

Same aproach as the other one \[f(x) = 3 \sqrt{-x + 3}\] We need to know in which interval the derivative is the smallest \[f'(x) =-\frac{ 1.5 }{ \sqrt{-x + 3} } \] The domain of \(f(x)\) is \( (\infty, 3]\) So only B is a valid interval Didn't even need the derivative nor taking a limit

OpenStudy (cwrw238):

right